# Origin of Pulsed Radio Emission from Magnetars

### v1

@@ 3 Simulation Method and Setup
Similar to the problem of radio emission in pulsars (Philippov et al., 2020), the superluminal mode is of special interest since, unlike the Alfvén mode, it does not suffer from Landau damping and can escape through a plasma of decreasing density $n$. As the superluminal mode travels through the plasma with decreasing $\omega_{\rm p}\propto\sqrt{n}$, it gradually becomes a usual electromagnetic wave with propagation speed approaching $c$ when $\omega_{\rm p}\ll\omega$. In our simulations, these modes propagate through the plasma and escape through the absorbing layers near the $y$-boundaries (see also Appendix E). At low $k$, the superluminal dispersion curve starts at the relativistic plasma frequency $\omega_{\rm min}=\sqrt{2}\omega_{\rm p}/\gamma_{\rm peak}^{3/2}$, where $\gamma_{\rm peak}=1/\sqrt{1-v_{\rm peak}^{2}}$. Its energy density observed inside the turbulence excitation region is concentrated around the local $\omega_{\rm min}$ (see Fig. 4). As a result, the characteristic frequency of escaping radiation is close to $\omega_{\rm min}$ defined inside the current-carrying region of the simulation box: $\omega_{\rm em}\approx\omega_{\rm min}$.
@@
Figure 5: Dependence of the normalized Poynting vector, $S_{\rm EM}$, of electromagnetic radiation escaping through the $y$-boundaries of 2X1V simulations on the cooling strength (panel a, for a fixed $v_{d}/c=0.5$), and the drift speed (panel b). Blue, orange and green data points in both panels correspond to simulations with cooling times $\omega_{p}t_{\rm cool}=300,1000,3000$. Errorbars represent fluctuations of the measured Poynting flux.
@@
The energy flux of escaping electromagnetic waves $S_{\rm EM}$ is described by the $y$-component of the Poynting vector $S_{y}=-cE_{x}\times B_{z}/4\pi$ measured in the boundary vacuum zones. We define $S_{\rm EM}\equiv(\int_{x}S_{y}{\rm{d}}x)/L_{x}$ by averaging $S_{y}$ along the $y$-boundaries of the simulation box. We find that for sufficiently large boxes $S_{\rm EM}$ shows only weak dependence on the box size. This indicates a saturated value for the energy density of superluminal waves, ${\cal E}_{\rm EM}\approx S_{\rm EM}/c$. We interpret this saturation as a balance established by the nonlinear interactions of plasma modes, partitioning the turbulence energy ${\cal E}_{\rm turb}$ between different modes, where ${\cal E}_{\rm turb}$ is a fraction of the total kinetic energy density of the plasma, ${\cal E}_{K}$. The energy density of superluminal waves then may be expected to scale with the total energy density of electrostatic waves obtained in 1D simulations, $\sim{\cal E}_{K}/(\omega_{p}t_{\rm cool})$.
@@
For our $v_{d}$ parameter scan, we also find the approximate consistency of the steady-state $S_{\rm EM}$ with Eq. (15) (panel b of Figure 5). The nonlinear saturation has a strong limiting effect on $S_{\rm EM}$ at $v_{d}\gtrsim 0.5$. In the model with $v_{d}=0.8c$, the observed $S_{\rm EM}$ falls below the analytical expression that uses ${\cal E}_{K}$ in the midplane $y=0$. This model has so strong turbulence that the zone of nonlinear saturation of the superluminal wave density ${\cal E}_{\rm EM}$ extends from the midplane to larger $y$ where our setup has a weaker electric current and hence a lower ${\cal E}_{K}$. The observed $S_{\rm EM}$ may become consistent with Eq. (15) when the appropriately reduced ${\cal E}_{K}$ is used in the analytical expression. We leave further study of this effect for future work.
For all simulations presented in this work, we also performed parallel simulations of the same setup using the PIC code Tristan-v2 (Hakobyan et al., 2023). We were unable to converge the results of PIC simulations for the amplitude of the weak initial electric field, $E_{0}$, and the peak amplitude of the coherent emission for our longest cooling time simulation, $t_{\rm cool}=3000\omega^{-1}_{p}$. Despite this limitation, PIC simulations gave converged results for the steady-state emission rate, which closely matches with our Gkeyll simulations.
@@ 6 A global model of radio emission
where $\nu_{p}=\sqrt{4\pi ne^{2}/m_{e}}=\gamma_{\star}^{1/2}\nu_{p}^{\prime}$ (since $n=\gamma_{\star}n^{\prime}$), and the plasma density $n$ is given by Eq. (3). As shown below, emission should occur in a plasma with $\gamma_{\star}\gg 1$ and a mildly relativistic $\gamma_{\rm peak}\sim 2$ in order to explain observations. Then, one can use the simplified estimate $\nu_{\rm em}\approx\gamma^{1/2}_{\star}\nu_{p}$. The values of $\nu_{\rm em}$ in our fiducial global magnetospheric model are shown in Figure 6.
@@
$S_{\rm EM}=\frac{\gamma_{\star}^{3}c\,eB\psi}{4\pi r_{e}R_{\max}}\frac{\alpha_{v^{\prime}}}{\omega^{\prime}_{p}t^{\prime}_{\rm cool}},$ (18)
where $r_{e}=e^{2}/m_{e}c^{2}$ is the classical electron radius. The apparent luminosity of radio emission (isotropic equivalent) in the lab frame is $L_{\rm EM}\approx 4\pi r^{2}S_{\rm EM}$, and we find
@@
As the radio waves propagate to the outer magnetosphere at $r>10^{3}R_{*}$, they can undergo cyclotron absorption. It occurs when the wave frequency transformed to the particle rest frame matches the cyclotron frequency, $\gamma(\omega-{\boldsymbol{k}}\cdot{\boldsymbol{v}})=\omega_{B}$. Cyclotron absorption is a significant problem for the escape of emission in ordinary radio pulsars (e.g., Philippov & Kramer, 2022). In Appendix F, we show that in the case of magnetars, the resonance occurs at $r_{\rm abs}\sim{\rm several}\times 10^{3}R_{*}$. We also find the absorption optical depth $\tau_{\rm abs}\sim(1-\cos\theta_{c}){\cal M}\sin^{2}\theta_{c}$ up to geometric factors of order unity (see Appendix F for the full expression), where $\theta_{c}$ is the latitude of the resonance location. The optical depth is reduced near the axis of the twisted magnetosphere because its current density $j_{0}$ (Eq. 2) and the corresponding plasma density (Eq. 3) drop near the axis. Radio waves produced at smaller radii will escape through the outer magnetosphere if they encounter $\tau_{\rm abs}<1$. This condition defines a polar “visibility zone” inside the emission region that produces escaping radiation. Its boundary is indicated by the blue curve in Figure 6. As one can see, the region of bright radio emission (generated with $\gamma_{\star}\sim 10$) lies within the visibility zone, and therefore it is not blocked by cyclotron absorption at $r_{\rm abs}$.
Note that the values of $\gamma_{\star}\sim 10$ are found in the proposed region of radio emission when radiative drag is dominated by X-rays flowing radially from the magnetar, as assumed in Figure 1. A more advanced model involves additional drag created by X-rays scattered in the magnetosphere, which reduces $\gamma_{\star}$ (Beloborodov, 2013). The contribution of scattered radiation depends on the anisotropy of magnetar’s surface emission (it may be dominated by the heated footprint of the twisted “$j$-bundle” around the magnetic pole) and the distribution of magnetospheric electric currents, which are gradually erased in the inner magnetosphere (Beloborodov, 2009). The complicated global picture will be further studied in future work.
@@ 7 Summary and Discussion
The radiative drag force serves as an engine that persistently drives the instability generating GHz radio waves. Therefore, the location of the radio-emitting region and the observed spectrum are determined by the properties of the radiation field in magnetar magnetospheres. Taking into account only thermal X-ray photons propagating radially from the magnetar, we find an extended emission region near the magnetic axis at radial distances ranging from 15 to 50 magnetar radii. This finding is consistent with the observed wide pulse profiles of radio emission. Future detailed calculations of the observed beam shape should take into account refraction, which is significant near the emission point where $\nu_{\rm em}\approx\nu_{p}$ (e.g., Lyubarskii & Petrova, 1998).
@@
The highest frequency in the radio-emitting zone, as calculated from a global two-fluid model, falls somewhat short of the $100$ GHz that is observed in radio magnetars (Torne et al., 2015). One possible explanation for this discrepancy is our simplistic assumption (adopted in the global model) that the observed emission frequency is tightly coupled to the local plasma frequency. In more detailed models, the frequency will be increased: in the simulations performed in large periodic boxes, which allow substantially longer times for nonlinear interactions, we find that the spectrum of superluminal modes extends to at least $\approx 3\nu^{\prime}_{p}$, pushing the predicted frequency cutoff towards $\approx 100\,{\rm GHz}$. We expect a similar increase of $\nu^{\prime}_{\rm em}$ for our simulations with non-periodic boundaries (which allow the escape of electromagnetic waves) when they are extended to larger simulation boxes. We will explore the mechanism of frequency increase in large boxes in future work.
@@
Polarization of radio emission has been observed in the canonical radio magnetar J1810–197 following outbursts in 2003 (Kramer et al., 2007) and 2018 (Dai et al., 2019), as well as in magnetars J1550–5418 (Camilo et al., 2008) and J1818–1607 (Fisher et al., 2024). Strongly linearly polarized radiation was detected in all cases, consistent with the production of coherent radiation in the highly magnetized pair plasma in the magnetosphere. Key polarization characteristics, including the swing of the position angle of the linearly polarized component and the degree of circular polarization as a function of the pulse phase, exhibit diverse and time-evolving characteristics. Although not always resembling the S-shape behavior predicted by the classic “rotating vector model” (RVM), in most cases, the position angle shows a significant swing that is consistent with the origin in the polar flux tube described in our model. The wave normal modes in the inner magnetosphere are linearly polarized. Thus, the often observed significant circular polarization (e.g., $\approx 15\%$ in J1550–5418 and 1818–1607 to nearly $50\%$ in parts of the profile of J1810–197 after the 2018 outburst) has to originate from propagation effects in the outer magnetosphere. The observed linear polarization of escaping radiation is also set at a large radius $r_{\rm pl}$ (e.g., Cheng & Ruderman, 1979; Petrova & Lyubarskii, 2000; Beskin & Philippov, 2012). Inside this radius, the difference of refractive indices of the two modes $\Delta n$ is sufficiently large to keep the wave in the local eigen mode. Hence, its polarization vector tracks the gradually changing direction of the local magnetic field $\boldsymbol{B}$, which may be described by angle $\phi_{B}$ in the plane perpendicular to the direction of propagation. The adiabatic tracking condition breaks at $r\sim r_{\rm pl}$ where $\Delta n\sim(\partial\phi_{B}/\partial r)c/\omega_{\rm em}$, and at larger radii the wave polarization vector “freezes out,” i.e. stops evolving along the ray. Evaluating this condition for typical parameters of plasma in magnetar magnetospheres gives $r_{\rm pl}/R_{*}\sim 10^{3}-10^{4}$, comparable to the cyclotron resonance radius. [^11] Explicit equations for $\Delta n$ and other characteristics of normal modes in the limiting polarization region are described in Appendix G. The high altitude of the polarization-limiting radius is consistent with significant deviations from RVM behavior and the appearance of substantial circular polarization (Petrova & Lyubarskii, 2000; Beskin & Philippov, 2012), although more careful calculations utilizing a realistic spatial plasma density distribution are required to make quantitative statements.
@@
We thank Roger Blandford, Ashley Bransgrove, Sasha Chernoglazov, Robbie Ewart, Hayk Hakobyan, Jens Mahlmann, Anatoly Spitkovsky, and Chris Thompson for fruitful discussions. This work was supported by NSF grant No. AST-2307395 (SZ and AP), Simons Foundation (00001470, AP), and facilitated by Multimessenger Plasma Physics Center (MPPC, AP and AB), NSF grant No. PHY-2206607. A.P. additionally acknowledges support by an Alfred P. Sloan Fellowship, and a Packard Foundation Fellowship in Science and Engineering. Computing resources were provided and supported by the Division of Information Technology at the University of Maryland (Zaratan cluster [^12] http://hpcc.umd.edu). This research is part of the Frontera computing project (Stanzione et al., 2020) at the Texas Advanced Computing Center (LRAC-AST21006). Frontera is made possible by NSF award OAC-1818253. This research was additionally supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@ Appendix A Two-stream Instability threshold for Lorentzian distributions
Criteria for the stability of hot counter-streaming $e^{\pm}$ beams can be computed exactly for the Lorentzian non-relativistic distribution function (e.g., O’Neil & Malmberg, 1968), which reads

### v2

@@ 3 Simulation Method and Setup
Similar to the problem of radio emission in pulsars (Philippov_2020), the superluminal mode is of special interest since, unlike the Alfvén mode, it does not suffer from Landau damping and can escape through a plasma of decreasing density $n$. As the superluminal mode travels through the plasma with decreasing $\omega_{\rm p}\propto\sqrt{n}$, it gradually becomes a usual electromagnetic wave with propagation speed approaching $c$ when $\omega_{\rm p}\ll\omega$. In our simulations, these modes propagate through the plasma and escape through the absorbing layers near the $y$-boundaries (see also Appendix E). At low $k$, the superluminal dispersion curve starts at the relativistic plasma frequency $\omega_{\rm min}=\sqrt{2}\omega_{\rm p}/\gamma_{\rm peak}^{3/2}$, where $\gamma_{\rm peak}=1/\sqrt{1-v_{\rm peak}^{2}}$. Its energy density observed inside the turbulence excitation region is concentrated around the local $\omega_{\rm min}$ (see Fig. 4). As a result, the characteristic frequency of escaping radiation is close to $\omega_{\rm min}$ defined inside the current-carrying region of the simulation box: $\omega_{\rm em}\approx\omega_{\rm min}$.
@@
Figure 5: Dependence of the normalized Poynting vector, $S_{\rm EM}$, of electromagnetic radiation escaping through the $y$-boundaries of 2X1V simulations on the cooling strength (panel (a), for a fixed $v_{d}/c=0.5$), and the drift speed (panel b). Blue, orange and green data points in both panels correspond to simulations with cooling times $\omega_{p}t_{\rm cool}=300,1000,3000$. The values represented by the dashed lines in panel (a) and the points in panel (b) are calculated from the data taken over the last $0.5t_{\rm cool}$ of each simulation. Errorbars represent fluctuations of the measured Poynting flux.
@@
The energy flux of escaping electromagnetic waves $S_{\rm EM}$ is described by the $y$-component of the Poynting vector $S_{y}=-cE_{x}\times B_{z}/4\pi$ measured in the boundary vacuum zones. [^11] In simulations conducted in periodic boxes, we find that the energy flux of superluminal waves is essentially isotropic in the RL frame. Consequently, we expect $S_{y}$ to be representative of the flux in any direction. We define $S_{\rm EM}\equiv(\int_{x}S_{y}{\rm{d}}x)/L_{x}$ by averaging $S_{y}$ along the $y$-boundaries of the simulation box. We find that for sufficiently large boxes $S_{\rm EM}$ shows only weak dependence on the box size. This indicates a saturated value for the energy density of superluminal waves, ${\cal E}_{\rm EM}\approx S_{\rm EM}/c$. We interpret this saturation as a balance established by the nonlinear interactions of plasma modes, partitioning the turbulence energy ${\cal E}_{\rm turb}$ between different modes, where ${\cal E}_{\rm turb}$ is a fraction of the total kinetic energy density of the plasma, ${\cal E}_{K}$. The energy density of superluminal waves then may be expected to scale with the total energy density of electrostatic waves obtained in 1D simulations, $\sim{\cal E}_{K}/(\omega_{p}t_{\rm cool})$.
@@
For our $v_{d}$ parameter scan, we also find the approximate consistency of the steady-state $S_{\rm EM}$ with Eq. (17) (panel (b) of Figure 5). The nonlinear saturation has a strong limiting effect on $S_{\rm EM}$ at $v_{d}\gtrsim 0.5$. In the model with $v_{d}=0.8c$, the observed $S_{\rm EM}$ falls below the analytical expression that uses ${\cal E}_{K}$ in the midplane $y=0$. This model has so strong turbulence that the zone of nonlinear saturation of the superluminal wave density ${\cal E}_{\rm EM}$ extends from the midplane to larger $y$ where our setup has a weaker electric current and hence a lower ${\cal E}_{K}$. The observed $S_{\rm EM}$ may become consistent with Eq. (17) when the appropriately reduced ${\cal E}_{K}$ is used in the analytical expression. We leave further study of this effect for future work.
For all simulations presented in this work, we also performed parallel simulations of the same setup using the PIC code Tristan-v2 (hayk_hakobyan_2023). We were unable to converge the results of PIC simulations for the amplitude of the weak initial electric field, $E_{0}$, and the peak amplitude of the coherent emission for our longest cooling time simulation, $t_{\rm cool}=3000\omega^{-1}_{p}$. Despite this limitation, PIC simulations gave converged results for the steady-state emission rate, which closely matches with our Gkeyll simulations.
@@ 6 A global model of radio emission
where $\nu_{p}=\sqrt{4\pi ne^{2}/m_{e}}/2\pi=\gamma_{\star}^{1/2}\nu_{p}^{\prime}$ (since $n=\gamma_{\star}n^{\prime}$), and the plasma density $n$ is given by Eq. (3). As shown below, emission should occur in a plasma with $\gamma_{\star}\gg 1$ and a mildly relativistic $\gamma_{\rm peak}\sim 2$ in order to explain observations. Then, one can use the simplified estimate $\nu_{\rm em}\approx\gamma^{1/2}_{\star}\nu_{p}$. The values of $\nu_{\rm em}$ in our fiducial global magnetospheric model are shown in Figure 6.
@@
$S_{\rm EM}=\frac{\gamma_{\star}^{3}c\,eB\psi}{4\pi r_{e}R_{\max}}\frac{\alpha_{v^{\prime}}}{\omega^{\prime}_{p}t^{\prime}_{\rm cool}}.$ (20)
The apparent luminosity of radio emission (isotropic equivalent) in the lab frame is $L_{\rm EM}\approx 4\pi r^{2}S_{\rm EM}$, and we find
@@
As the radio waves propagate to the outer magnetosphere at $r>10^{3}R_{*}$, they can undergo cyclotron absorption. It occurs when the wave frequency transformed to the particle rest frame matches the cyclotron frequency, $\gamma(\omega-{\boldsymbol{k}}\cdot{\boldsymbol{v}})=\omega_{B}$. Cyclotron absorption is a significant problem for the escape of emission in ordinary radio pulsars (e.g., Philippov_2022). In Appendix F, we show that in the case of magnetars, the resonance occurs at $r_{\rm abs}\sim{\rm several}\times 10^{3}R_{*}$. We also find the absorption optical depth $\tau_{\rm abs}\sim(1-\cos\theta_{c}){\cal M}\sin^{2}\theta_{c}$ up to geometric factors of order unity (see Appendix F for the full expression), where $\theta_{c}$ is the latitude of the resonance location. The optical depth is reduced near the axis of the twisted magnetosphere because its current density $j_{0}$ (Eq. 2) and the corresponding plasma density (Eq. 3) drop near the axis. Radio waves produced at smaller radii will escape through the outer magnetosphere if they encounter $\tau_{\rm abs}<1$. This condition defines a polar “visibility zone” inside the emission region that produces escaping radiation. Its boundary is indicated by the blue curve in Figure 6. As one can see, the region of bright radio emission (generated with $\gamma_{\star}\sim 10$) lies within the visibility zone, and therefore it is not blocked by cyclotron absorption at $r_{\rm abs}$.
Note that the values of $\gamma_{\star}\sim 10$ are found in the proposed region of radio emission when radiative drag is dominated by X-rays flowing radially from the magnetar, as assumed in Figure 1. A more advanced model involves additional drag created by X-rays scattered in the magnetosphere, which reduces $\gamma_{\star}$ (Beloborodov_2013). The contribution of scattered radiation depends on the anisotropy of magnetar’s surface emission (it may be dominated by the heated footprint of the twisted “$j$-bundle” around the magnetic pole) and the distribution of magnetospheric electric currents, which are gradually erased in the inner magnetosphere (Beloborodov_2009). The complicated global picture will be further studied in future work.
@@ 7 Summary and Discussion
The radiative drag force serves as an engine that persistently drives the instability generating GHz radio waves. Therefore, the location of the radio-emitting region and the observed spectrum are determined by the properties of the radiation field in magnetar magnetospheres. Taking into account only thermal X-ray photons propagating radially from the magnetar, we find an extended emission region near the magnetic axis at radial distances ranging from 15 to 50 magnetar radii. This finding is consistent with the observed wide pulse profiles of radio emission. Future detailed calculations of the observed beam shape should take into account refraction, which is significant near the emission point where $\nu_{\rm em}\approx\nu_{p}$ (e.g., LyubPet98).
@@
The highest frequency in the radio-emitting zone, as calculated from a global two-fluid model, falls somewhat short of the $100$ GHz that is observed in radio magnetars (Torne2015). One possible explanation for this discrepancy is our simplistic assumption (adopted in the global model) that the observed emission frequency is tightly coupled to the local plasma frequency. In more detailed models, the frequency will be increased: in the simulations performed in large periodic boxes, which allow substantially longer times for nonlinear interactions, we find that the spectrum of superluminal modes extends to at least $\approx 3\nu^{\prime}_{p}$, pushing the predicted frequency cutoff towards $\approx 100\,{\rm GHz}$. We expect a similar increase of $\nu^{\prime}_{\rm em}$ for our simulations with non-periodic boundaries (which allow the escape of electromagnetic waves) when they are extended to larger simulation boxes. We will explore the mechanism of frequency increase in large boxes in future work.
@@
Polarization of radio emission has been observed in the canonical radio magnetar J1810–197 following outbursts in 2003 (KramerStappers2007) and 2018 (Dai2019), as well as in magnetars J1550–5418 (Camilo_2008) and J1818–1607 (Fisher2024). Strongly linearly polarized radiation was detected in all cases, consistent with the production of coherent radiation in the highly magnetized pair plasma in the magnetosphere. Key polarization characteristics, including the swing of the position angle of the linearly polarized component and the degree of circular polarization as a function of the pulse phase, exhibit diverse and time-evolving characteristics. Although not always resembling the S-shape behavior predicted by the classic “rotating vector model” (RVM), in most cases, the position angle shows a significant swing that is consistent with the origin in the polar flux tube described in our model. The wave normal modes in the inner magnetosphere are linearly polarized. Thus, the often observed significant circular polarization (e.g., $\approx 15\%$ in J1550–5418 and 1818–1607 to nearly $50\%$ in parts of the profile of J1810–197 after the 2018 outburst) has to originate from propagation effects in the outer magnetosphere. The observed linear polarization of escaping radiation is also set at a large radius $r_{\rm pl}$ (e.g., ChengRuderman1979; PetLyub2000; BeskinPh2012). Inside this radius, the difference of refractive indices of the two modes $\Delta n$ is sufficiently large to keep the wave in the local eigen mode. Hence, its polarization vector tracks the gradually changing direction of the local magnetic field $\boldsymbol{B}$, which may be described by angle $\phi_{B}$ in the plane perpendicular to the direction of propagation. The adiabatic tracking condition breaks at $r\sim r_{\rm pl}$ where $\Delta n\sim(\partial\phi_{B}/\partial r)c/\omega_{\rm em}$, and at larger radii the wave polarization vector “freezes out,” i.e. stops evolving along the ray. Evaluating this condition for typical parameters of plasma in magnetar magnetospheres gives $r_{\rm pl}/R_{*}\sim 10^{3}-10^{4}$, comparable to the cyclotron resonance radius. [^12] Explicit equations for $\Delta n$ and other characteristics of normal modes in the limiting polarization region are described in Appendix G. The high altitude of the polarization-limiting radius is consistent with significant deviations from RVM behavior and the appearance of substantial circular polarization (PetLyub2000; BeskinPh2012), although more careful calculations utilizing a realistic spatial plasma density distribution are required to make quantitative statements.
@@
We thank Roger Blandford, Ashley Bransgrove, Sasha Chernoglazov, Robbie Ewart, Hayk Hakobyan, Jens Mahlmann, Anatoly Spitkovsky, and Chris Thompson for fruitful discussions. This work was supported by NSF grant No. AST-2307395 (SZ and AP), Simons Foundation (00001470, AP), and facilitated by Multimessenger Plasma Physics Center (MPPC, AP and AB), NSF grant No. PHY-2206607. A.P. additionally acknowledges support by an Alfred P. Sloan Fellowship, and a Packard Foundation Fellowship in Science and Engineering. Computing resources were provided and supported by the Division of Information Technology at the University of Maryland (Zaratan cluster [^13] http://hpcc.umd.edu). This research is part of the Frontera computing project (Frontera) at the Texas Advanced Computing Center (LRAC-AST21006). Frontera is made possible by NSF award OAC-1818253. This research was additionally supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@ Appendix A Two-stream Instability threshold for Lorentzian distributions
Criteria for the stability of hot counter-streaming $e^{\pm}$ beams can be computed exactly for the Lorentzian non-relativistic distribution function (e.g., ONeil_1968), which reads